Randy needed to climb a ladder to get on the roof of his house. He leaned his 15 foot ladder against the edge of his roof, which is 10 feet off the ground. How far must the ladder be away from the building in order for it to touch the side of his house?
1
Expert's answer
2012-10-30T10:16:01-0400
we use Pythagorean theorem. hypotenuse=15 one of the leg10 so the another leg x
x^2+10^2=15^2
x=sqrt(15^2-10^2)=sqrt(125)=5sqrt(5)
so the answer 5sqrt(5)